OpenStudy (anonymous):

if excess Na3 P04 reacts with 650.0cm^3 of 1.25M FeCl2 solution,what mass of Fe3(PO4)2 will be made

OpenStudy (anonymous):

\[2Na _{3}PO _{4}+3FeCl _{2}\rightarrow Fe _{3}(PO _{4})_{2}+6NaCl\]

OpenStudy (espex):

So 3 moles of FeCl2 is \[(3 \times 55.85) + (2 \times 35.45)=238.45g\]

OpenStudy (anonymous):

would that be my answer?

OpenStudy (espex):

No, it is part of my thought process. What you are looking for is this, you have a 3:1 ratio of FeCl2:Fe3(PO4)2, you are given a solution and need to figure out how many moles of it you have and then use the ratio to evaluate the amount they say you have.

OpenStudy (espex):

So basically your limiting factor for this reaction is the FeCl2, now the part that I have difficulty with is the conversions from molarity to mass.

OpenStudy (espex):

650 cm^3 is equal to 650ml. \[650mL \times \frac{1.25moles}{1000 mL} = 0.8125 moles\]

OpenStudy (espex):

\[\frac{3mole}{1mole}=\frac{0.8125mole}{xmol} \rightarrow 3x=0.8125 \rightarrow x=0.271mol\]

OpenStudy (espex):

\[\frac{3mol FeCl_2}{6Fe_3(PO_4)_2}=\frac{0.8125mol FeCl_2}{x mol Fe_3(PO_4)_2} = 3x=4.875 \rightarrow x=1.625 mol Fe_3(PO_4)_2\]

OpenStudy (espex):

(3*55.85)+(2(30.97+4*16.00)=357.49g

OpenStudy (espex):

\[\frac{357.49g}{1mol} \times 1.625mol=580.9g\]

OpenStudy (anonymous):

okay i think i get it thanks so much lol i have to go now im distracteed with te internet and i need to finish a worksheet

OpenStudy (espex):

:)

OpenStudy (anonymous):

thanks lol i was tourturing myself on this Q

OpenStudy (espex):

You're welcome, it looks like it's right to me, but I would also suggest you compare it to your text. :) Good luck.

OpenStudy (anonymous):

thanks lol we dont use textbooks we use whatever the teacher gives us on notes thanks for the good luck too ima need it for my tests tomorrow